# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 4.1.1 Single person 3D pose estimation in images
Solving Depth Ambiguity. As shown in Fig.4(a), different 3D pose coordinates projecting to 2D images may give the same results, leading to an ill-posed problem. This problem can be solved by using the propagation properties of light and camera imaging principle. To address the ill-posed problem, Wei et al. wei2019view introduced a view-invariant framework to moderate the effects of viewpoint diversity. A View-Invariant Hierarchical Correction (VI-HC) network predicts the 3D pose refinement with view-consistent constraints in the proposed framework. Additionally, a view-invariant discriminative network actualizes high-level constraints after the base network generates an initial estimation. Ray-based 3D (Ray3D) absolute estimation method zhan2022ray3d can convert the input from pixel space to 3D normalized rays. Another helpful approach for this problem is learning a more appropriate feature representation. Part-centric HEatMap triplets (HEMlets) framework zhou2021hemlets utilizes three joint-heatmaps to represent the end-joints relative depth information, which bridges the gap between the 2D location and the 3D human pose.
@@
Solving Body Structure Understanding. Unlike other computer vision tasks, the body’s unique structures can provide constraints or prior information to improve pose estimation performance. In the Joint Relationship Aware Network (JRAN) zheng2020joint , a dual attention module was designed to generate both whole and local feature attention block weights. The limb poses aware network wu2021limb leverages kinematic constraint and trajectory information to prevent errors from accumulating along the human body structure. Pose grammar in xu2021monocular learns a 2D-3D mapping function, enabling the input 2D pose to be transformed into 3D space, and integrates three aspects of human structure (kinematics, symmetry, motor coordination) into a Bi-directional Recurrent Neural Network (RNN). The orientation keypoints-based method fisch2021orientation uses virtual markers to generate sufficient information for accurately inferring rotations through simple post-processing. The Graph Neural Network (GNN) is a convolutional network defined on the graph data structure. It is difficult for typical neural networks to handle graph-structured data, such as body structure information. However, this challenge can be more easily addressed if graph-based learning methods are used, as shown in Fig.4(b). Liu et al. liu2019feature proposed a feature boosting network in which features are learned by the convolutional layers and are boosted with Graphical ConvLSTM to perceive the graphical long short-term dependency among different body parts. The Locally Connected Network (LCN) ci2020locally leverages the allocation of dedicated filters rather than sharing them for different joints. Modulated Graph Convolutional Network (Modulated-GCN) zou2021modulated includes weight and affinity modulation to learn modulation vectors for different body joints and to adjust the graph structure, respectively. Zeng et al. zeng2021learning designed a skeletal GNN learning framework to address the depth ambiguity and self-occlusion problems in 3D human pose estimation. In this framework, the proposed hop-aware hierarchical channel-squeezing fusion layer is designed to extract relevant information from neighboring nodes. Higher-order Regular Splitting graph Network (RS-Net) hassan2023regular captures long-range dependencies between body joints using multi-hop neighborhoods. It learns distinct modulation vectors for different body joints and adds modulation matrices to the corresponding skeletal adjacency matrices. Zhai et al. zhai2023hopfir employed intra-group joint refinement utilizing attention mechanisms to discover potential joint synergies to explore the potential synergies between joints.
@@
Solving Occlusion Problems. Partial occlusion of the human body, including self-occlusion and other occlusions (such as object occlusion and multi-person occlusion), is typical in various scenes. Occlusion may interfere with pose estimation, potentially resulting in the prediction of an error pose. Solving the occlusion problem through a multi-view approach is effective, where an occluded pose not visible in one view is likely to be visible in other views in a multi-view system. Thus, a multi-view based approach can provide more reliable results through cross-view frame inference. In practice, Iskakov et al. iskakov2019learnable combined algebraic and volumetric triangulation for 3D pose estimation from multi-view 2D images. The former method is based on a differentiable algebraic triangulation with confidence weights, while the latter method utilizes volumetric aggregation from intermediate 2D backbone feature maps. Multi-view geometric priors have been incorporated in qiu2019cross , which combines two steps: first, predicting the 2D poses in multi-view RGB images through cross-view fusion, and second, utilizing the proposed Recursive Pictorial Structure Model (RPSM) to recover the 3D poses from the previously predicted multi-view 2D poses. To improve the real-time performance of multi-view 3D pose estimation, Remelli et al. remelli2020lightweight designed a lightweight framework with a differentiable Direct Linear Transform (DLT) layer. Adaptive multi-view fusion zhang2021adafuse enhances the features in occluded views by determining the correspondence between points in occluded and visible views. To address the problem of generalizability for multi-view estimation, Bartol et al. bartol2022generalizable proposed a stochastic learning framework for pose triangulation. Luvizon et al. luvizon2022consensus effectively integrated 2D annotated data and 3D poses to design a consensus-aware method, which optimizes multi-view poses from uncalibrated images by different coherent estimations up to a scale factor from the intrinsic parameters. Probabilistic triangulation module jiang2023probabilistic embedded in a 3D pose estimation framework can extend multi-view methods to uncalibrated scenes. It models camera poses using probability distributions and iteratively updates them from 2D features, replacing the traditional update through camera poses and eliminating the dependency on camera calibration.
@@
Unsupervised learning typically focuses on learning features rather than specific tasks from unlabeled training data, and it finds relationships between samples by mining the intrinsic features of the data. The first work kudo2018unsupervised can predict 3D human pose without any 3D dataset by adversarial learning, which is based on the generative adversarial networks via unsupervised training. The approach chen2019unsupervised utilizes geometric self-supervision and randomly reprojects 2D pose camera viewpoints from the recovered 3D skeleton during training, thereby forming a self-consistency loss in the lift-reproject-lift process. Elepose wandt2022elepose utilizes random projections, estimating likelihood using normalizing flows on 2D poses in a linear subspace. Furthermore, Elepose also learns the distribution of camera angles to reduce the dependence on camera rotation priors in the training dataset.
@@
Self-supervised learning is a branch of unsupervised learning where the model can learn by itself from unlabeled training data and acquire the representation model on unlabeled data through pretext tasks. During on-the-ground execution, EpipolarPose kocabas2019self trains the 3D pose estimator without any 3D ground-truth data or camera extrinsic parameters, predicting 2D poses from multi-view images and obtaining a 3D pose and camera geometry via epipolar geometry. Wang et al. wang20193d designed a simple yet effective self-supervised correction mechanism for 3D human pose estimation by learning all intrinsic structures of the human pose, which is divided into two parts: the 2D-to-3D pose transformation and 3D-to-2D pose projection. MRP-Net kundu2022uncertainty reformulated the self-supervised 3D human pose estimation as an unsupervised domain adaptation problem, which includes model-free joint localization and model-based parametric regression. To reduce dependence on the consistency loss that guides learning, unlike previous works, the method gong2022posetriplet generates 2D-3D pose pairs for augmenting supervision via the proposed self-enhancing dual-loop learning framework.
@@
A weakly-supervised model, characterized by its reliance on only weak labels, often faces the challenge of performing complex tasks with limited or imprecise guidance. Despite the lack of detailed annotations typically required in fully supervised scenarios, these models derive meaningful insights, utilizing vague or less informative labels to infer intricate patterns and relationships within the data. This ability to operate with minimal supervision makes them particularly useful when acquiring comprehensive labeled data is impractical or costly. In practical applications, Hua et al. hua2022weakly proposed a weakly-supervised method that initially lifts the 2D keypoints into coarse 3D poses across two views using triangulation and then refines the 3D pose by employing spatial configurations and cross-view correlations. Yang et al. yang2023camerapose designed a weakly-supervised framework that utilizes projection relationships to estimate 3D poses solely from 2D pose annotations under the condition of known camera parameters. Transfer learning enables the transfer of model parameters from the source domain to the target domain, allowing us to share learned model parameters with a new model, as shown in Fig.4(c). This approach speeds up and optimizes efficiency, eliminating the need to learn from scratch, as is common with most networks. For example, Adaptpose gholami2022adaptpose is an end-to-end cross-dataset adaptation 3D human pose estimation predictor that is implemented using transfer learning for limited data applications.
@@ 4.1.2 Single person 3D pose estimation in videos
With hardware development, image data are often acquired and processed in videos. The video frame data can provide more continuous information than a single-frame image, predicting human pose estimation in sequence through the spatio-temporal domain. In addition, optical flow and scene flow information can be extracted from videos to predict 3D human pose from data of multimodal tang2023ftcm ; zhao2023poseformerv2 . Solving Single-frame Limitation. Using continuous image frames from video can capture dynamic changes and temporal information. By analyzing this information, deep learning models can extract motion features and spatio-temporal relationships, thereby effectively improving the effectiveness of 3D human pose estimation. Pavllo et al. pavllo20193d proposed a 3D human pose estimation method for video that extracts temporal cues with dilated convolutions over 2D keypoint trajectories, and then they designed a semi-supervised method with back-projection to improve performance. PoseFormer zheng20213d as a spatial-temporal transformer-based framework predicts 3D body pose by analyzing human joint relations within each frame and their temporal correlations across frames. Based on previous work artacho2020unipose , unipose+ artacho2021uniposeplus leverages multi-scale feature representations to enhance the feature extractors in the framework’s backbone. This framework utilizes contextual information across scales and joint localization with Gaussian heatmap modulation to improve decoder estimation accuracy. Furthermore, Multi-Hypothesis transFormer (MHFormer) li2022mhformer learns spatio-temporal representations with multiple plausible pose hypotheses in a three-stage framework. In Mixed Spatio-Temporal Encoder (MixSTE) framework zhang2022mixste , the temporal transformer block learns the temporal motion of each joint and their inter-joint spatial correlation. Honari et al. honari2022temporal proposed an unsupervised feature extraction method using Contrastive Self-Supervised (CSS) learning to extract temporal information from videos. Extending this line of research, Hierarchical Spatial-Temporaltrans Formers (HSTFormer) qian2023hstformer utilizes spatial-temporal correlations of joints at different levels simultaneously, marking a first in studying hierarchical transformer encoders with multi-level fusion. Recently, Spatio-Temporal Criss-cross (STC) attention block tang20233d decomposes correlation learning into space and time to reduce the computational cost of the joint-to-joint affinity matrix.
@@
Solving Real-time Problems. However, while various methods improve 3D pose estimation performance, they also introduce new problems and challenges in video-based estimation. For example, the continuous video frames add a substantial computational cost, which poses a challenge to the processing efficiency of the system and makes real-time processing more difficult. To reduce the total computational complexity, Einfalt et al. einfalt2023uplift designed a transformer-based scheme that uplifts dense 3D poses from temporally sparse 2D pose sequences by temporal upsampling within Transformer blocks. Sun et al. sun2023mixsynthformer reduced the complexity of attention calculation in Transformers through spatio-temporal sparse sampling, enabling the estimation of 3D human poses in video sequences on computationally constrained platforms.
@@
Solving Body Structure Understanding. At the same time, the continuous motion information provided by video has led researchers to combine videos with human kinematics. Wang et al. wang2020motion designed a motion loss that computes the difference between the motion patterns of the predicted and ground truth keypoint trajectories, along with motion encoding, a simple yet effective representation of keypoint motion for 3D human pose estimation in videos. Dynamical Graph Network (DG-Net) zhang2021learning dynamically determines the human-joint affinity and adaptively predicts human pose via spatial-temporal joint relations in videos. To address the shortcoming of directly regressing the 3D joint locations, Chen et al. chen2021anatomy proposed an anatomy-aware estimation framework. This human skeleton anatomy-based framework includes a bone direction prediction network and a bone length prediction network, and it effectively utilizes bone features to bridge the gap between 2D keypoints and 3D joints. Xue et al. xue2022boosting designed a part-aware temporal attention module capable of extracting each part’s temporal dependency separately.
@@
Solving Occlusion Problems. To address occlusion problems, video-based 3D human pose estimation enables the use of a multi-view approach and leverages the continuity of inputs between video frames to predict occluded body parts. Cheng et al. cheng2019occlusion introduced an occlusion-aware model that utilizes estimated 2D confidence heatmaps of keypoints and an optical-flow consistency constraint to generate a more complete 3D pose in occlusion scenes. Additionally, Multi-view and Temporal Fusing Transformer (MTF-Transformer) shuai2022adaptive fuses multi-view sequences in uncalibrated scenes for 3D human pose estimation in videos. To reduce dependency on camera calibration, the framework infers the relationship between pairs of views with a relative attention mechanism.
@@
Solving Data Lacking. Human actions are inherently continuous, making video-based pose estimation critical for enhanced understanding. Compared with 3D annotated still images, high-quality, annotated 3D videos are relatively scarce. Nevertheless, the Internet abounds with a vast repository of unlabeled video data, the utilization of which for learning purposes holds considerable importance. Thus, video-based 3D human pose estimation to alleviate data dependency is essential. In practice, Yu et al. yu2021towards divided the unsupervised 3D pose estimation process into two sub-tasks: a scale estimation module and a pose lifting module. The scale estimation module optimizes the 2D input pose, while the pose lifting module maps the optimized 2D pose to its 3D counterpart. In weakly-supervised methods, Chen et al. chen2019weakly proposed a weakly-supervised method, employing a view synthesis framework and a geometry-aware representation. In this framework, the method leverages 2D keypoints for supervision and learns a shared 3D representation between viewpoints by synthesizing the human pose from one viewpoint to another. Semi-supervised learning utilizes only partially labeled data, consisting of a large amount of unlabeled data and a small amount of labeled data. In this context, Mitra et al. mitra2020multiview present a multi-view consistent semi-supervised method, which can regress 3D human pose from unannotated, uncalibrated, but synchronized multi-view videos. In self-supervised methods, Kundu et al. kundu2020self employed prior knowledge of the body skeleton and disentangles the inherent factors of variation through part-guided human image synthesis. Similarly, Shan et al. shan2022p randomly mask the body joints in spatial and temporal domains to better capture spatial and temporal dependencies. Meta-learning (a.k.a. ”learning to learn”) is a process where algorithms optimize their learning strategy based on experience. It involves training a model on various learning tasks, enabling it to learn new tasks more efficiently using fewer data samples. Cho et al. cho2021camera present an optimization-based meta-learning algorithm for 3D human pose estimation. This algorithm can adapt to arbitrary camera distortion by generating synthetic distorted data from undistorted 2D keypoints during model training. Apart from changing the learning supervision methods, data augmentation is also a practical approach to increasing the amount of data and enhancing the model’s generalization performance. PoseAug gong2021poseaug adjusts geometric factors such as posture, body shape, and viewpoint for model learning under the premise of the discriminative module controlling the feasibility of augmented poses. Zhang et al. zhang2023learning generates more diverse and challenging poses online.

### v2

@@ 4.1.1 Single person 3D pose estimation in images
Solving Depth Ambiguity. As shown in Fig.5(a), different 3D pose coordinates projecting to 2D images may give the same results, leading to an ill-posed problem. This problem can be solved by using the propagation properties of light and camera imaging principle. To address the ill-posed problem, Wei et al. [63] introduced a view-invariant framework to moderate the effects of viewpoint diversity. A View-Invariant Hierarchical Correction (VI-HC) network predicts the 3D pose refinement with view-consistent constraints in the proposed framework. Additionally, a view-invariant discriminative network actualizes high-level constraints after the base network generates an initial estimation. Ray-based 3D (Ray3D) absolute estimation method [64] can convert the input from pixel space to 3D normalized rays. Another helpful approach for this problem is learning a more appropriate feature representation. Part-centric HEatMap triplets (HEMlets) framework [65] utilizes three joint-heatmaps to represent the end-joints relative depth information, which bridges the gap between the 2D location and the 3D human pose.
@@
Solving Body Structure Understanding. Unlike other computer vision tasks, the body’s unique structures can provide constraints or prior information to improve pose estimation performance. In the Joint Relationship Aware Network (JRAN) [66], a dual attention module was designed to generate both whole and local feature attention block weights. The limb poses aware network [67] leverages kinematic constraint and trajectory information to prevent errors from accumulating along the human body structure. Pose grammar in [68] learns a 2D-3D mapping function, enabling the input 2D pose to be transformed into 3D space, and integrates three aspects of human structure (kinematics, symmetry, motor coordination) into a Bi-directional Recurrent Neural Network (RNN). The orientation keypoints-based method [69] uses virtual markers to generate sufficient information for accurately inferring rotations through simple post-processing.
The Graph Neural Network (GNN) is a convolutional network defined on the graph data structure. It is difficult for typical neural networks to handle graph-structured data, such as body structure information. However, this challenge can be more easily addressed if graph-based learning methods are used, as shown in Fig.5(b). Liu et al. [70] proposed a feature boosting network in which features are learned by the convolutional layers and are boosted with Graphical ConvLSTM to perceive the graphical long short-term dependency among different body parts. The Locally Connected Network (LCN) [71] leverages the allocation of dedicated filters rather than sharing them for different joints. Modulated Graph Convolutional Network (Modulated-GCN) [72] includes weight and affinity modulation to learn modulation vectors for different body joints and to adjust the graph structure, respectively. Zeng et al. [73] designed a skeletal GNN learning framework to address the depth ambiguity and self-occlusion problems in 3D human pose estimation. In this framework, the proposed hop-aware hierarchical channel-squeezing fusion layer is designed to extract relevant information from neighboring nodes. Higher-order Regular Splitting graph Network (RS-Net) [59] captures long-range dependencies between body joints using multi-hop neighborhoods. It learns distinct modulation vectors for different body joints and adds modulation matrices to the corresponding skeletal adjacency matrices. Zhai et al. [74] employed intra-group joint refinement utilizing attention mechanisms to discover potential joint synergies to explore the potential synergies between joints.
@@
Solving Occlusion Problems. Partial occlusion of the human body, including self-occlusion and other occlusions (such as object occlusion and multi-person occlusion), is typical in various scenes. Occlusion may interfere with pose estimation, potentially resulting in the prediction of an error pose. Solving the occlusion problem through a multi-view approach is effective, where an occluded pose not visible in one view is likely to be visible in other views in a multi-view system. Thus, a multi-view based approach can provide more reliable results through cross-view frame inference. In practice, Iskakov et al. [75] combined algebraic and volumetric triangulation for 3D pose estimation from multi-view 2D images. The former method is based on a differentiable algebraic triangulation with confidence weights, while the latter method utilizes volumetric aggregation from intermediate 2D backbone feature maps. Multi-view geometric priors have been incorporated in [76], which combines two steps: first, predicting the 2D poses in multi-view RGB images through cross-view fusion, and second, utilizing the proposed Recursive Pictorial Structure Model (RPSM) to recover the 3D poses from the previously predicted multi-view 2D poses. To improve the real-time performance of multi-view 3D pose estimation, Remelli et al. [77] designed a lightweight framework with a differentiable Direct Linear Transform (DLT) layer. Adaptive multi-view fusion [78] enhances the features in occluded views by determining the correspondence between points in occluded and visible views. To address the problem of generalizability for multi-view estimation, Bartol et al. [79] proposed a stochastic learning framework for pose triangulation. Luvizon et al. [80] effectively integrated 2D annotated data and 3D poses to design a consensus-aware method, which optimizes multi-view poses from uncalibrated images by different coherent estimations up to a scale factor from the intrinsic parameters. Probabilistic triangulation module [31] embedded in a 3D pose estimation framework can extend multi-view methods to uncalibrated scenes. It models camera poses using probability distributions and iteratively updates them from 2D features, replacing the traditional update through camera poses and eliminating the dependency on camera calibration.
@@
Unsupervised learning typically focuses on learning features rather than specific tasks from unlabeled training data, and it finds relationships between samples by mining the intrinsic features of the data. The first work [81] can predict 3D human pose without any 3D dataset by adversarial learning, which is based on the generative adversarial networks via unsupervised training. The approach [82] utilizes geometric self-supervision and randomly reprojects 2D pose camera viewpoints from the recovered 3D skeleton during training, thereby forming a self-consistency loss in the lift-reproject-lift process. Elepose [83] utilizes random projections, estimating likelihood using normalizing flows on 2D poses in a linear subspace. Furthermore, Elepose also learns the distribution of camera angles to reduce the dependence on camera rotation priors in the training dataset.
@@
Self-supervised learning is a branch of unsupervised learning where the model can learn by itself from unlabeled training data and acquire the representation model on unlabeled data through pretext tasks. During on-the-ground execution, EpipolarPose [84] trains the 3D pose estimator without any 3D ground-truth data or camera extrinsic parameters, predicting 2D poses from multi-view images and obtaining a 3D pose and camera geometry via epipolar geometry. Wang et al. [85] designed a simple yet effective self-supervised correction mechanism for 3D human pose estimation by learning all intrinsic structures of the human pose, which is divided into two parts: the 2D-to-3D pose transformation and 3D-to-2D pose projection. MRP-Net [86] reformulated the self-supervised 3D human pose estimation as an unsupervised domain adaptation problem, which includes model-free joint localization and model-based parametric regression. To reduce dependence on the consistency loss that guides learning, unlike previous works, the method [58] generates 2D-3D pose pairs for augmenting supervision via the proposed self-enhancing dual-loop learning framework.
@@
A weakly-supervised model, characterized by its reliance on only weak labels, often faces the challenge of performing complex tasks with limited or imprecise guidance. Despite the lack of detailed annotations typically required in fully supervised scenarios, these models derive meaningful insights, utilizing vague or less informative labels to infer intricate patterns and relationships within the data. This ability to operate with minimal supervision makes them particularly useful when acquiring comprehensive labeled data is impractical or costly. In practical applications, Hua et al. [87] proposed a weakly-supervised method that initially lifts the 2D keypoints into coarse 3D poses across two views using triangulation and then refines the 3D pose by employing spatial configurations and cross-view correlations. Yang et al. [47] designed a weakly-supervised framework that utilizes projection relationships to estimate 3D poses solely from 2D pose annotations under the condition of known camera parameters.
Transfer learning enables the transfer of model parameters from the source domain to the target domain, allowing us to share learned model parameters with a new model, as shown in Fig.5(c). This approach speeds up and optimizes efficiency, eliminating the need to learn from scratch, as is common with most networks. For example, Adaptpose [88] is an end-to-end cross-dataset adaptation 3D human pose estimation predictor that is implemented using transfer learning for limited data applications.
@@ 4.1.2 Single person 3D pose estimation in videos
With hardware development, image data are often acquired and processed in videos. The video frame data can provide more continuous information than a single-frame image, predicting human pose estimation in sequence through the spatio-temporal domain. In addition, optical flow and scene flow information can be extracted from videos to predict 3D human pose from data of multimodal [122, 60].
@@
Solving Single-frame Limitation. Using continuous image frames from video can capture dynamic changes and temporal information. By analyzing this information, deep learning models can extract motion features and spatio-temporal relationships, thereby effectively improving the effectiveness of 3D human pose estimation. Pavllo et al. [89] proposed a 3D human pose estimation method for video that extracts temporal cues with dilated convolutions over 2D keypoint trajectories, and then they designed a semi-supervised method with back-projection to improve performance. PoseFormer [90] as a spatial-temporal transformer-based framework predicts 3D body pose by analyzing human joint relations within each frame and their temporal correlations across frames. Based on previous work [123], unipose+ [91] leverages multi-scale feature representations to enhance the feature extractors in the framework’s backbone. This framework utilizes contextual information across scales and joint localization with Gaussian heatmap modulation to improve decoder estimation accuracy.
Furthermore, Multi-Hypothesis transFormer (MHFormer) [92] learns spatio-temporal representations with multiple plausible pose hypotheses in a three-stage framework. In Mixed Spatio-Temporal Encoder (MixSTE) framework [93], the temporal transformer block learns the temporal motion of each joint and their inter-joint spatial correlation. Honari et al. [94] proposed an unsupervised feature extraction method using Contrastive Self-Supervised (CSS) learning to extract temporal information from videos. Extending this line of research, Hierarchical Spatial-Temporaltrans Formers (HSTFormer) [95] utilizes spatial-temporal correlations of joints at different levels simultaneously, marking a first in studying hierarchical transformer encoders with multi-level fusion. Recently, Spatio-Temporal Criss-cross (STC) attention block [96] decomposes correlation learning into space and time to reduce the computational cost of the joint-to-joint affinity matrix.
@@
Solving Real-time Problems. However, while various methods improve 3D pose estimation performance, they also introduce new problems and challenges in video-based estimation. For example, the continuous video frames add a substantial computational cost, which poses a challenge to the processing efficiency of the system and makes real-time processing more difficult. To reduce the total computational complexity, Einfalt et al. [21] designed a transformer-based scheme that uplifts dense 3D poses from temporally sparse 2D pose sequences by temporal upsampling within Transformer blocks. Sun et al. [97] reduced the complexity of attention calculation in Transformers through spatio-temporal sparse sampling, enabling the estimation of 3D human poses in video sequences on computationally constrained platforms.
@@
Solving Body Structure Understanding. At the same time, the continuous motion information provided by video has led researchers to combine videos with human kinematics. Wang et al. [98] designed a motion loss that computes the difference between the motion patterns of the predicted and ground truth keypoint trajectories, along with motion encoding, a simple yet effective representation of keypoint motion for 3D human pose estimation in videos. Dynamical Graph Network (DG-Net) [99] dynamically determines the human-joint affinity and adaptively predicts human pose via spatial-temporal joint relations in videos. To address the shortcoming of directly regressing the 3D joint locations, Chen et al. [100] proposed an anatomy-aware estimation framework. This human skeleton anatomy-based framework includes a bone direction prediction network and a bone length prediction network, and it effectively utilizes bone features to bridge the gap between 2D keypoints and 3D joints. Xue et al. [101] designed a part-aware temporal attention module capable of extracting each part’s temporal dependency separately.
@@
Solving Occlusion Problems. To address occlusion problems, video-based 3D human pose estimation enables the use of a multi-view approach and leverages the continuity of inputs between video frames to predict occluded body parts. Cheng et al. [102] introduced an occlusion-aware model that utilizes estimated 2D confidence heatmaps of keypoints and an optical-flow consistency constraint to generate a more complete 3D pose in occlusion scenes. Additionally, Multi-view and Temporal Fusing Transformer (MTF-Transformer) [32] fuses multi-view sequences in uncalibrated scenes for 3D human pose estimation in videos. To reduce dependency on camera calibration, the framework infers the relationship between pairs of views with a relative attention mechanism.
@@
Solving Data Lacking. Human actions are inherently continuous, making video-based pose estimation critical for enhanced understanding. Compared with 3D annotated still images, high-quality, annotated 3D videos are relatively scarce. Nevertheless, the Internet abounds with a vast repository of unlabeled video data, the utilization of which for learning purposes holds considerable importance. Thus, video-based 3D human pose estimation to alleviate data dependency is essential. In practice, Yu et al. [103] divided the unsupervised 3D pose estimation process into two sub-tasks: a scale estimation module and a pose lifting module. The scale estimation module optimizes the 2D input pose, while the pose lifting module maps the optimized 2D pose to its 3D counterpart. In weakly-supervised methods, Chen et al. [104] proposed a weakly-supervised method, employing a view synthesis framework and a geometry-aware representation. In this framework, the method leverages 2D keypoints for supervision and learns a shared 3D representation between viewpoints by synthesizing the human pose from one viewpoint to another. Semi-supervised learning utilizes only partially labeled data, consisting of a large amount of unlabeled data and a small amount of labeled data. In this context, Mitra et al. [105] present a multi-view consistent semi-supervised method, which can regress 3D human pose from unannotated, uncalibrated, but synchronized multi-view videos. In self-supervised methods, Kundu et al. [106] employed prior knowledge of the body skeleton and disentangles the inherent factors of variation through part-guided human image synthesis. Similarly, Shan et al. [107] randomly mask the body joints in spatial and temporal domains to better capture spatial and temporal dependencies. Meta-learning (a.k.a. ”learning to learn”) is a process where algorithms optimize their learning strategy based on experience. It involves training a model on various learning tasks, enabling it to learn new tasks more efficiently using fewer data samples. Cho et al. [57] present an optimization-based meta-learning algorithm for 3D human pose estimation. This algorithm can adapt to arbitrary camera distortion by generating synthetic distorted data from undistorted 2D keypoints during model training. Apart from changing the learning supervision methods, data augmentation is also a practical approach to increasing the amount of data and enhancing the model’s generalization performance. PoseAug [108] adjusts geometric factors such as posture, body shape, and viewpoint for model learning under the premise of the discriminative module controlling the feasibility of augmented poses. Zhang et al. [109] generates more diverse and challenging poses online.